# Computing an approximation of the partial Weyl closure of a holonomic module

### v1

@@ 6.1. Gröbner bases
For the above examples, the parenthesis (gl) in Table 1 indicates that the order
@@
was used in Julia and Singular, while the weight vector $(\mathbf{1},\mathbf{1})$ was chosen in Macaulay2. Similarly, the parenthesis (elim) indicates that the block order
@@
This last example is taken from (Levandovskyy and Morales, 2008). A star (*) after the timing of a computation with Macaulay2 indicates that the returned Gröbner basis is different than the other two due to the use of a different monomial order.
@@
Gröbner basis over $\mathbb{Q}$
@@ 6.2. Weyl closure
where on the last equation
@@
The symbols $\operatorname{J},\operatorname{I},\operatorname{Y},\operatorname{K}$ denote the usual Bessel functions and $\operatorname{Ai}$ denotes the Airy function of the first kind. Eqs. 28 and 30 were taken from (Chyzak, 2014), Eq. 29 appeared in (Beukers, 1983), Eq. 33 belongs to a family of integrals related to the enumeration of small step walks in the quarter plane (Bostan et al., 2017) (with a polynomial factor in the numerator removed), Eq. 36 is from the database (Koutschan, 2014), and Eq. 37 is the period of a Calabi–Yau 3-fold appearing in the database associated to the article (Batyrev and Kreuzer, 2010) (Topology #42, polytope v25.59). The functions Eqs. 31 and 32 and Eqs. 35 and 34 are respectively obtained from Eq. 30 and Eq. 33 by artificially increasing the severity of their singularities. This is achieved by multiplying the original functions by a factor of the form $\exp(1/u)$ or $\log(u)$, where $u$ is a polynomial vanishing on their corresponding singular locus. Note that this benchmark table does not contain any case of partial Weyl closure since there are no algorithms to compare against.

### v2

@@ 6.1. Gröbner bases
For the above examples, the label (gl) in Table 1 indicates that the order
@@
was used in Julia and Singular, while the weight vector $(\mathbf{1},\mathbf{1})$ was chosen in Macaulay2. Similarly, the label (elim) indicates that the block order
@@
This last example is taken from (Levandovskyy and Morales, 2008). A star (*) after the timing of a computation with Macaulay2 indicates that the returned Gröbner basis is different than the other two due to the use of a different monomial order. MEM indicates that the computation ran out of memory.
@@
Gröbner bases over $\mathbb{Q}$
@@ 6.2. Weyl closure
where in the last equation
@@
The symbols $\operatorname{J},\operatorname{I},\operatorname{Y},\operatorname{K}$ denote the usual Bessel functions and $\operatorname{Ai}$ denotes the Airy function of the first kind. Eqs. 28 and 30 were taken from (Chyzak, 2014), Eq. 29 appeared in (Beukers, 1983), Eq. 33 belongs to a family of integrals related to the enumeration of small step walks in the quarter plane (Bostan et al., 2017) (with a polynomial factor in the numerator removed), Eq. 36 is from the database (Koutschan, 2014), and Eq. 37 is the period of a Calabi–Yau 3-fold appearing in the database associated with the article (Batyrev and Kreuzer, 2010) (Topology #42, polytope v25.59). The functions Eqs. 31, 32, 35 and 34 are obtained from Eqs. 30 and 33 by artificially increasing the severity of their singularities. This is achieved by multiplying the original functions by a factor of the form $\exp(1/u)$ or $\log(u)$, where $u$ is a polynomial vanishing on their corresponding singular locus. Note that this benchmark table does not include any case of partial Weyl closure since there are no algorithms to compare against.